OpenStudy (anonymous):

Derivate: f(x)=ln(sqrt(x) + sqrt(x+1)) f'(1) = ? I don't really understand wolframs steps. So could anyone help me :(.

OpenStudy (anonymous):

...

OpenStudy (lgbasallote):

chain rule... first let x = (sqrt x + sqrt x +1 so 1/x(dx) dx is the derivative of x which is sqrt x + sqrt x+1 take the derivative of that... 1/2sqrtx + 1/2sqrt(x+1) so your answer is (1/2sqrtx + 1/2sqrtx+1)/x simplify the numerator... 1/2sqrtx + 1/2sqrtx+1 LCD = 2(sqrtx)(sqrtx+1) so sqrt(x+1) + sqrtx/2(sqrtx)(sqrtx+1)/x simplify... sqrt(x+1) + sqrtx/2x(sqrtx)(sqrtx+1) was that what wolframalpha gave you?

OpenStudy (lgbasallote):

oh...forgot to sub... sqrt2 + 1/2sqrt2

OpenStudy (anonymous):

wolfram says: 1/(2 sqrt(x) sqrt(1+x))

OpenStudy (lgbasallote):

oh i'm sorry!!! i missubstituted...i made the denominator x when it was supposed to be sqrtx + sqrtx+1...lemme redo that...

OpenStudy (mimi_x3):

\[\large \log(\sqrt{x} +\sqrt{x+1}) \] Use: \[\frac{d}{dx} \log(..) = \left(\frac{1}{(..)}\right) * (..)'\] When you find the derivative of: \[\sqrt{x} +\sqrt{x+1} \] =>\[\frac{1}{2}\left(\frac{1}{\sqrt{x+1}}+\frac{1}{\sqrt{x}}\right) \] \[\frac{1}{\sqrt{x}+\sqrt{x+1}} *\frac{1}{2}\left(\frac{1}{\sqrt{x+1}}+\frac{1}{\sqrt{x}}\right) \] Maybe..It looks like it might match Wolfram Alpha's Answer.

OpenStudy (anonymous):

@mimi, it's not log, it's natural logarithm. .

OpenStudy (mimi_x3):

loolol, typo.

OpenStudy (mimi_x3):

but its still the same answer i suppose.

OpenStudy (anonymous):